AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 15 of 33  ELOG logo
ID Date Author Subject Text Attachments
  379   Mon Oct 3 16:03:33 2022 OH, CBTTT14 PhotosPhotos of the new TTT14 DSSD delivered to
GSI.
Photos of detectors are broken up by photos
  
  378   Sun Aug 14 06:50:58 2022 TD, JMSunday 14 August
07.51 DAQ continues files 22Feb/R97
 10x 
  377   Sat Aug 13 10:23:59 2022 JM, TDNoise tests pt3 and pulser walkWaveforms and discriminator ON. Pulser
re-connected to all fees via daisy chain.
V=100v, I=8.36uA
 15x 
  376   Fri Aug 12 15:12:37 2022 JM, TDCARME measurementsmeasurement 1 -string
intermediate (smaller) chamber - 2144
(mm) circumference,  measured half circumference
 9x 
  375   Fri Aug 12 09:06:57 2022 TD, JMNoise tests pt2
10.07 DSSSD bias -100V leakage current -8.65uA
 30x 
  374   Thu Aug 11 09:25:13 2022 JM, TDNoise tests
Moved pneumatic motor in and out several
revolutions
 14x 
  373   Wed Aug 10 10:33:03 2022 TD, ML, JM, NPMeeting at YR09 to discuss options for CARME operations to February 2022
Initial discussions of proposals to provide
open CARME chamber and provide access to
 20220810_111638.jpg20220810_095036.jpg20220810_095150.jpg20220810_095159.jpg 
  372   Wed Aug 10 10:30:04 2022 TD, JMWednesday 10 August
09.45 Found CARME electrical power off
 4.PNG3.PNG2.png1.png 
  371   Tue Aug 9 16:07:33 2022 JM, TDAnydesk setup command lineTo log into appc218 desktop:
same login and password as to carme@atppc022.gsi.de

To start anydesk
  
  370   Tue Aug 9 13:44:16 2022 TD, JMTuesday 9 August
14.00 CARME vacuum c. 1.8e-10 mbar - see
attachment 1
 Capture_1.PNG20220809_163450.jpg20220809_135840.jpg20220809_135833.jpg20220809_135801.jpg 
  369   Thu Mar 3 15:14:21 2022 JM, OHCARME packing upPulser walkthrough taken for motor position
23.1mm and motor position 67.43mm.

Points taken at 30,50,70,90,110,130,150,170,190,210,230
  
  368   Wed Mar 2 22:29:48 2022 CB, SDLast shift with beam22:30 Shift handed over. DAQ OK. Target
OK. Beam OK.

 
 9x 
  367   Wed Mar 2 21:16:06 2022 JM, PMCurrent status22:16

Temperature setpoint = 113k

Not observed noise burst again
  
  366   Wed Mar 2 19:25:25 2022 JM, PMCurrent status20:25

Observed short burst of low energy
noise on aida1, lasted <30s (attachment
 10x 
  365   Wed Mar 2 17:21:35 2022 JM, PMCurrent status18:21

Temperature setpoint = 113k

Target density OK S1 varies between
 8x 
  364   Wed Mar 2 14:15:16 2022 JM, PMCurrent status15:15

Temperature setpoint = 113 K

Target density is OK for now. S1
 6x 
  363   Wed Mar 2 09:38:50 2022 Spillmann, Uweion source restarted 

detector position

out: 23.1mm
  
  362   Wed Mar 2 06:00:49 2022 Spillmann, Uwewaiting07:00
Target density is hovering between
2.0e-8 and 2.3e-8
  
  361   Wed Mar 2 03:40:40 2022 David, GeorgeTemperature Cycle04:40
Target density is hovering between
2.0e-8 and 2.3e-8, sometimes dipping below
  
  360   Wed Mar 2 01:24:48 2022 David, GeorgeIon Source Update02:25
Zoran was unable to ignite the ion
source.
  
ELOG V3.1.4-unknown